OpenStudy (anonymous):

need help rationalizing the denominator \[2ab/\sqrt{5}-1\]

OpenStudy (anonymous):

Multiply the numerator and denominator by \[\sqrt{5}\]

OpenStudy (anonymous):

so i end up with \[2ab \sqrt{5}/4\]?

OpenStudy (anonymous):

Is it \[ \frac{ 2ab }{ \sqrt{5} -1} ?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Know about conjugates?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

With surds: two surds are said to be conjugate of each other, if their product gives a difference of their square

OpenStudy (anonymous):

So in this case we're multiplying the denominator by it's conjugate, to rationalize it...

OpenStudy (anonymous):

Conjugate of \[(\sqrt{5}-1) = (\sqrt{5}+1)\]

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good, then work it out...

OpenStudy (anonymous):

so i should have an answer of 2ab(root5=1) / 6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

No, try again.

OpenStudy (anonymous):

i dont understand how to multiply the \[\sqrt{5}-1\]and \[\sqrt{5}+1\]

OpenStudy (anonymous):

You should get\[\frac{ 2ab (\sqrt{5} +1)}{ 4}\] Reduced to: \[\frac{ab(\sqrt{5} +1)}{ 2}\]

OpenStudy (anonymous):

oh okay so the only thing i was doing wrong was not reducing the fraction?

OpenStudy (anonymous):

Okay, multiplying that is trivial. Just expand as before

OpenStudy (anonymous):

i see now. so to multiply that i just cancel the root making 5 just 5 and then subtract the 1

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

okay thank you very much!

OpenStudy (anonymous):

Like this \[(\sqrt{5}-1)(\sqrt{5}+1)\] \[5 +\sqrt{5}-\sqrt{5}-1=4\]

OpenStudy (anonymous):

You have to pay special attention to the signs. You miss, that everything gets messed up!

OpenStudy (anonymous):

Yw. (:

OpenStudy (anonymous):

:) great help!!

OpenStudy (anonymous):

Can you attempt this? \[\frac{ 4-8\sqrt{2} }{ 1+\sqrt{2}}\]